The volume of a metal sphere increases by 0.15% when its temperature is raised by 24°C. The coefficient of linear expansion of metal is?

Grade:12th pass

1 Answers

Yashraj Singh
7 years ago
Let € be coefficient of volumetric expansionLet coefficient of linear expansion be $Vnew=V(1+€∆t)V+0.15% of V=V(1+€20)1V+0.0015V=V(1+€20)1.0015V=V(1+€20)0.0015/20=€3/8*10^-4=€3$=€3$=3/8*10^-4$=1.25*10^-5